# On the Worst Prompt Performance of Large Language Models

### v2

@@ Abstract
The performance of large language models (LLMs) is acutely sensitive to the phrasing of prompts, which raises significant concerns about their reliability in real-world scenarios. Existing studies often divide prompts into task-level instructions and case-level inputs and primarily focus on evaluating and improving robustness against variations in tasks-level instructions. However, this setup fails to fully address the diversity of real-world user queries and assumes the existence of task-specific datasets. To address these limitations, we introduce RobustAlpacaEval, a new benchmark that consists of semantically equivalent case-level queries and emphasizes the importance of using the worst prompt performance to gauge the lower bound of model performance. Extensive experiments on RobustAlpacaEval with ChatGPT and six open-source LLMs from the Llama, Mistral, and Gemma families uncover substantial variability in model performance; for instance, a difference of 45.48% between the worst and best performance for the Llama-2-70B-chat model, with its worst performance dipping as low as 9.38%. We further illustrate the difficulty in identifying the worst prompt from both model-agnostic and model-dependent perspectives, emphasizing the absence of a shortcut to characterize the worst prompt. We also attempt to enhance the worst prompt performance using existing prompt engineering and prompt consistency methods, but find that their impact is limited. These findings underscore the need to create more resilient LLMs that can maintain high performance across diverse prompts. Data and code are available at https://github.com/cbwbuaa/On-the-Worst-Prompt-Performance-of-LLMs.
@@ 1 Introduction
In recent years, large language models (LLMs) have made extraordinary progress 1, 2, 3, 4 and one key factor in their success is their ability to adapt to diverse tasks through prompting. However, the performance of these LLMs is significantly sensitive to the prompts they receive 5, 6, 7. Even minor alterations in format, without semantic changes, can trigger substantial performance degradation 8. Consequently, prompt engineering has surfaced as a critical component, playing a critical role in unlocking the full potential of these models 9, 10, 11, 12, 13.
Despite the efficacy of prompt engineering, it is not without its drawbacks. First, it is unrealistic to expect users to master the art of designing optimal prompts or to invest a significant amount of time in doing so. Second, automatic prompt engineering often necessitates testing on substantial labeled data to pinpoint the most effective prompts 14, 12, a process that is impractical for users who just have one or a few unlabeled queries. Therefore, investigating LLMs’ robustness to prompt variances is of great research and practical value.
@@
Previous research on prompt robustness 6, 8, 7 divides a user query into two parts: task-level instruction and case-level input. The task-level instruction is an abstract task definition and explanation supplemented by case-level concrete input (See the left part of 1). This setup is reminiscent of conventional NLP practice, where task-specific models are developed using a testing set for each well-defined NLP task. Consequently, they have focused on the LLM’s resilience to task-level instructions exclusively, reporting the average performance across testing set cases. Nonetheless, these studies have limitations. First, they disregard that the best task-level instruction might vary across individual cases. Second, they overlook the impact of variations in case-level input on model performance. Last but not least, real-world user queries often do not explicitly segregate task-level instruction and case-level input (See the right part of 1). These queries may cover a wide array of tasks and it is not possible to optimize the prompts through evaluating on a task-specific testing set.
In this paper, we present a comprehensive study that goes beyond the conventional approach of evaluating LLMs. We shift the focus from task-level instructions to diverse real-world user queries. Our work introduces a new benchmark, RobustAlpacaEval, that includes semantically equivalent case-level queries across various tasks, offering a more holistic analysis. We argue that the worst prompt performance is a very important metric for assessing the lower bound of LLM performance. Our extensive experiments on ChatGPT and six open-source LLMs from the Llama, Mistral, and Gemma families reveal substantial variability in model performance. For instance, Llama-2-70B-chat model shows a difference of up to 45.48 points in win-rate against GPT4 using RobustAlpacaEval, and its worst prompt performance can be as low as 9.38%. Our findings further reveal that the worst prompts, which lead to notably poor model performance, are not universally applicable across different models. Each model also exhibits unique preferences towards all paraphrases, as demonstrated by their inconsistent performance rankings on identical cases. Beyond the absence of model-agnostic traits in identifying the worst prompts, model-dependent prompt features, such as perplexity and hidden states, also prove inadequate in forecasting model performance. Furthermore, we also find that different models are more sensitive in different cases, which further underscores the complexity involved in tackling the issue of worst prompt performance.
@@ 2 Related Work
Previous studies 5, 15, 16, 17 have explored the robustness of LLMs to intentional perturbations such as word deletion and sentence shuffling. Our work deviates from these perturbation-based studies, as we are interested in model performance across semantically equivalent and syntactically fluent prompts. While it is anticipated that a flawed prompt would lead to performance decline, we do not expect such performance variations in our setting.
Recent research 6, 8, 7 has also examined the variability in model performance with semantically equivalent prompts. However, these studies solely focused on task-level instructions and overlooked the variations within case-level inputs. Moreover, the models used in these studies are trained and evaluated on traditional NLP datasets, while today’s LLMs are predominantly instruction-tuned on and serve diverse user prompts. This discrepancy suggests that the findings of these studies may not fully apply to the latest generation of LLMs.
@@
A number of studies utilize gradient-based methods for prompt optimization 9, 14, 18, 19, 20, 21. However, their dependence on substantial labeled training data and gradient computing limits their applicability for users with only a few unlabeled queries or gradient-less APIs. Furthermore, these methods incur significant computational costs as the model scale increases.
Conversely, gradient-free methods aim to find the optimal prompt through exploration and scoring 12, 10, 11, 22, 23. During the exploration phase, they create analogous prompts using techniques such as rephrasing the current prompts with an LLM. Subsequently, the best prompts are selected based on the model’s performance on downstream tasks. This process can be iterative. Although these approaches can proficiently pinpoint prompts that result in better performance for a specific task, they all necessitate a test set for prompt scoring. However, in real-world scenarios, such test sets are often absent. Gonen et al. 13 suggests that prompt perplexity can be an effective data-free criterion. However, our experiment results tell a different story, indicating that prompt perplexity is not a reliable metric.
@@ 3.1 A New Benchmark: RobustAlpacaEval
Our benchmark is based on TinyAlpaceEval 24, which is a condensed subset of the AlpacaEval 25 benchmark, created to enable efficient assessment of LLMs. We develop RobustAlpacaEval by creating ten paraphrases for each query within TinyAlpacaEval. To save manual efforts, this is first accomplished automatically through GPT4. Subsequently, each paraphrase is manually reviewed and revised to ensure semantic integrity and human-like fluency. We use carefully crafted prompts to promote the diversity of paraphrases, which is evidenced by the fact that the average of length-normalized edit distance [^1] We compute the edit distance between prompts $x$ and $y$ normalized by their average length as: $2*\text{EditDistance}(x,y)/(|x|+|y|)$ between each pair of paraphrases is 0.7234 at the word level. The specific few-shot paraphrasing instruction utilized in this process is detailed in Appendix A.
@@
In line with common practice, we use weighted win-rate 25 as our performance metric; It uses an evaluator to compare the output of the target model against that of a reference model, and estimates the winning probability of the target model. Specifically, we employ the gpt4_turbo model as the evaluator and the reference model. We term the model’s performance on the original prompt as original performance. We also report the worst, best, average performances across all eleven prompts as well as the standard deviation. For each metric, we average the results across all cases in RobustAlpacaEval.
@@ 3.2 Results
Model Kendall’s W Agreement Levels Distribution
Negligible Weak Moderate Strong
Llama family 0.443 0 0.242 0.414 0.343
Gemma family 0.548 0 0.08 0.28 0.64
Llama/Gemma/Mistral-7B 0.401 0.011 0.326 0.411 0.253
All models 0.238 0.053 0.723 0.202 0.021
@@ 4.1 Model-agnostic Analysis
We continue to check if the rankings among different models are consistent. We utilize Kendall’s W 26, commonly referred to as the coefficient of concordance, as a means to quantify the correlation among performance rankings derived from various models. Mathematically, Kendall’s W is defined as follows:
@@
where $R_{i}$ represents the sum of the ranks assigned to the $i$-th prompt by all models. $\bar{R}$ is the mean of total ranks, calculated as $\bar{R}=\frac{1}{n}\sum_{i=1}^{n}{R_{i}}$. $k$ is the number of models, and $n$ is the number of prompts being ranked. The value of W ranges from 0 to 1, where 0 indicates no agreement and 1 indicates complete agreement among the judges [^2] W values less than 0.1 are often considered as no or negligible agreement; 0.1 to 0.3 suggest weak agreement; 0.3 to 0.5 indicate moderate agreement; and W values above 0.5 show strong agreement 27..
@@ 4.2 Model-dependent Analysis
  • Prompt Perplexity. Gonen et al. 13 report that lower prompt perplexity generally correlates with superior model performance. To verify this assertion, we compute the Pearson correlation coefficients between perplexity and performance across all cases. The results are illustrated in Figure 3 (left half). The layout divides correlations into negative and positive (left/right of the 0 tick), with correlation strengths categorized as weak/no (0 to 0.3), moderate (0.3 to 0.6), and strong (0.6 to 1). Analyzing the Llama family models, we note the following: (i) About 45% of cases show weak or no correlation. (ii) Positive and negative correlations are nearly equally distributed. (iii) The distribution of cases across different correlation degrees is fairly balanced. These results are consistent with observations from other models analyzed (detailed in Appendix B), indicating no definitive correlation between prompt perplexity and model performance.
  • Min-KProb. We also explore the potential relationship between model performance and the prompt’s Min-KProb 28, a method related to pre-training data detection, which judges the model’s familiarity with the text based on the average log-likelihood of the lowest ktokens in the sentence. Intrigued by this, we carry out an analysis akin to that for perplexity, only to find weak positive correlations that lack statistical significance, as shown in the right half of Figure 3.
  • Hidden States. We shift our focus to investigate model perceptions of prompt quality through the lens of model representation, using: (i) Principal Component Analysis (PCA) for direct visualization of hidden states, as shown in Figure 5 (a). The color gradient reflects the model’s performance ranking across 11 prompts per case, revealing no distinct separation between different performance levels. (ii) Probing technique to quantify the extent to which the model’s internal representations are predictive of its performance. We train a reward model (a 3-layer MLP in practice) that scores prompts based on hidden states and apply the standard training loss function 29, which is defined as:
@@ 5 Improving Worst Prompt Performance
Drawing inspiration from the swarm distillation method proposed by Zhou et al. 30, we implement an unsupervised consistency regularization strategy that encourages the model’s predictions for various paraphrases to converge.
@@
    where $\bar{S_{ij}}$ is the mean of $S_{ij}$ over j. The value of $C(X,Y)$ ranges from 0 to $\infty$, reflecting the degree of consistency in the model’s log confidence of generating outputs ($Y$) based on given instructions ($X$), with 0 indicating totally consistent. We obverse that the consistency always decreases at first, followed by increasing, which aligns with the findings in Zhou et al. 30. We train models based on LoRA and stop training at the checkpoint from which $C(X,Y)$ starts to increase.
@@ 6 Conclusion
Appendix A Instruction for Paraphrasing

### v4

# Bowen Cao♢, Deng Cai♡, Zhisong Zhang Yuexian Zou Wai Lam

@@ Abstract
The performance of large language models (LLMs) is acutely sensitive to the phrasing of prompts, which raises significant concerns about their reliability in real-world scenarios. Existing studies often divide prompts into task-level instructions and case-level inputs and primarily focus on evaluating and improving robustness against variations in tasks-level instructions. However, this setup fails to fully address the diversity of real-world user queries and assumes the existence of task-specific datasets. To address these limitations, we introduce RobustAlpacaEval, a new benchmark that consists of semantically equivalent case-level queries and emphasizes the importance of using the worst prompt performance to gauge the lower bound of model performance. Extensive experiments on RobustAlpacaEval with ChatGPT and six open-source LLMs from the Llama, Mistral, and Gemma families uncover substantial variability in model performance; for instance, a difference of 45.48% between the worst and best performance for the Llama-2-70B-chat model, with its worst performance dipping as low as 9.38%. We further illustrate the difficulty in identifying the worst prompt from both model-agnostic and model-dependent perspectives, emphasizing the absence of a shortcut to characterize the worst prompt. We also attempt to enhance the worst prompt performance using existing prompt engineering and prompt consistency methods, but find that their impact is limited. These findings underscore the need to create more resilient LLMs that can maintain high performance across diverse prompts. Data and code are available at https://github.com/bwcao/RobustAlpacaEval.
@@ 1 Introduction
In recent years, large language models (LLMs) have made extraordinary progress 1, 2, 3, 4 and one key factor in their success is their ability to adapt to diverse tasks through prompting. However, the performance of these LLMs is significantly sensitive to the prompts they receive 5, 6, 7, 8. Even minor alterations in format, without semantic changes, can trigger substantial performance degradation 9. Consequently, prompt engineering has surfaced as a critical component, playing a critical role in unlocking the full potential of these models 10, 11, 12, 13, 14, 15.
Despite the efficacy of prompt engineering, it is not without its drawbacks. First, it is unrealistic to expect users to master the art of designing optimal prompts or to invest a significant amount of time in doing so. Second, automatic prompt engineering often necessitates testing on substantial labeled data to pinpoint the most effective prompts 16, 13, a process that is impractical for users who just have one or a few unlabeled queries. Therefore, investigating LLMs’ robustness to prompt variances is of great research and practical value.
@@
Previous research on prompt robustness 6, 9, 7 divides a user query into two parts: task-level instruction and case-level input. The task-level instruction is an abstract task definition and explanation supplemented by case-level concrete input (See the left part of Figure 1). This setup is reminiscent of conventional NLP practice, where task-specific models are developed using a testing set for each well-defined NLP task. Consequently, they have focused on the LLM’s resilience to task-level instructions exclusively, reporting the average performance across testing set cases. Nonetheless, these studies have limitations. First, they disregard that the best task-level instruction might vary across individual cases. Second, they overlook the impact of variations in case-level input on model performance. Last but not least, real-world user queries often do not explicitly segregate task-level instruction and case-level input (See the right part of Figure 1). These queries may cover a wide array of tasks and it is not possible to optimize the prompts through evaluating on a task-specific testing set.
In this paper, we present a comprehensive study that goes beyond the conventional approach of evaluating LLMs. We shift the focus from task-level instructions to diverse real-world user queries. Our work introduces a new benchmark, RobustAlpacaEval, that includes semantically equivalent case-level queries across various tasks, offering a more holistic analysis. We argue that the worst prompt performance, defined as the lowest performance a model exhibits across different paraphrases of a query with equal semantics and fluency, is a crucial metric for assessing the lower bound of LLM performance. Our extensive experiments on ChatGPT (gpt-3.5-turbo-1106) and six open-source LLMs from the Llama, Mistral, and Gemma families reveal substantial variability in model performance. For instance, Llama-2-70B-chat model shows a difference of up to 45.48 points in win-rate against GPT4 using RobustAlpacaEval, and its worst prompt performance can be as low as 9.38%. Our findings further reveal that the worst prompts, which cause the model to perform the worst in specific cases, are not universally applicable across different models. Each model also exhibits unique preferences towards all paraphrases, as demonstrated by their inconsistent performance rankings on identical cases. Beyond the absence of model-agnostic traits in identifying the worst prompts, model-dependent prompt features, such as perplexity and hidden states, also prove inadequate in forecasting model performance. Furthermore, we also find that different models are more sensitive in different cases, which further underscores the complexity involved in tackling the issue of worst prompt performance.
@@ 2 Related Work
Previous studies 5, 17, 18, 19, 20 have explored the robustness of LLMs to intentional perturbations such as word deletion and sentence shuffling. Our work deviates from these perturbation-based studies, as we are interested in model performance across semantically equivalent and syntactically fluent prompts. While it is anticipated that a flawed prompt would lead to performance decline, we do not expect such performance variations in our setting.
Recent research 6, 9, 7 has also examined the variability in model performance with semantically equivalent prompts. However, these studies solely focused on task-level instructions and overlooked the variations within case-level inputs. Moreover, the models used in these studies are trained and evaluated on traditional NLP datasets, while today’s LLMs are predominantly instruction-tuned on and serve diverse user prompts. This discrepancy suggests that the findings of these studies may not fully apply to the latest generation of LLMs.
@@
A number of studies utilize gradient-based methods for prompt optimization 10, 16, 21, 22, 23, 24. However, their dependence on substantial labeled training data and gradient computing limits their applicability for users with only a few unlabeled queries or gradient-less APIs. Furthermore, these methods incur significant computational costs as the model scale increases.
Conversely, gradient-free methods aim to find the optimal prompt through exploration and scoring 13, 11, 12, 25, 26. During the exploration phase, they create analogous prompts using techniques such as rephrasing the current prompts with an LLM. Subsequently, the best prompts are selected based on the model’s performance on downstream tasks. This process can be iterative. Although these approaches can proficiently pinpoint prompts that result in better performance for a specific task, they all necessitate a test set for prompt scoring. However, in real-world scenarios, such test sets are often absent. Gonen et al. 14 suggests that prompt perplexity can be an effective data-free criterion. However, our experiment results tell a different story, indicating that prompt perplexity is not a reliable metric.
@@ 3.1 A New Benchmark: RobustAlpacaEval
Our benchmark is based on TinyAlpaceEval 27, which is a condensed subset of the AlpacaEval 28 benchmark, created to enable efficient assessment of LLMs. We develop RobustAlpacaEval by creating ten [^1] We find that 10 paraphrases are sufficient to steadily reflect the variability in the model’s performance. Further details can be found in Appendix A.1. paraphrases for each query within TinyAlpacaEval. To save manual efforts, this is first accomplished automatically through GPT4. Subsequently, each paraphrase is manually reviewed and revised to ensure semantic integrity and human-like fluency. We use carefully crafted prompts to promote the diversity of paraphrases, which is evidenced by the fact that the average of length-normalized edit distance [^2] We compute the edit distance between prompts $x$ and $y$ normalized by their average length as: $2*\text{EditDistance}(x,y)/(|x|+|y|)$ between each pair of paraphrases is 0.7234 at the word level. The few-shot paraphrasing instruction utilized in this process is detailed in Appendix A.2.
@@
In line with common practice, we use weighted win-rate 28 as our performance metric; It uses an evaluator to compare the output of the target model against that of a reference model, and estimates the winning probability of the target model. Specifically, we employ the gpt4_turbo model as the evaluator and the reference model. We term the model’s performance on the original prompt as original performance. We also report the worst, best, average performances across all eleven prompts as well as the standard deviation. For each metric, we average the results across all cases in RobustAlpacaEval.
@@ 4.1 Model-agnostic Analysis
We continue to check if the rankings among different models are consistent. We utilize Kendall’s W 29, commonly referred to as the coefficient of concordance, as a means to quantify the correlation among performance rankings derived from various models. Mathematically, Kendall’s W is defined as follows:
@@
where $R_{i}$ represents the sum of the ranks assigned to the $i$-th prompt by all models. $\bar{R}$ is the mean of total ranks, calculated as $\bar{R}=\frac{1}{n}\sum_{i=1}^{n}{R_{i}}$. $k$ is the number of models, and $n$ is the number of prompts being ranked. The value of W ranges from 0 to 1, where 0 indicates no agreement and 1 indicates complete agreement among the judges [^3] W values less than 0.1 are often considered as no or negligible agreement; 0.1 to 0.3 suggest weak agreement; 0.3 to 0.5 indicate moderate agreement; and W values above 0.5 show strong agreement 30..
Model Kendall’s W Agreement Levels Distribution
Negligible Weak Moderate Strong
Llama family 0.443 0 0.242 0.414 0.343
Gemma family 0.548 0 0.08 0.28 0.64
Llama/Gemma/Mistral-7B 0.401 0.011 0.326 0.411 0.253
All models 0.238 0.053 0.723 0.202 0.021
@@ 4.2 Model-dependent Analysis
  • Prompt Perplexity. Gonen et al. 14 report that lower prompt perplexity generally correlates with superior model performance. To verify this assertion, we compute the Pearson correlation coefficients between perplexity and performance across all cases. The results are illustrated in Figure 4 (left half). The layout divides correlations into negative and positive (left/right of the 0 tick), with correlation strengths categorized as weak/no (0 to 0.3), moderate (0.3 to 0.6), and strong (0.6 to 1). Analyzing the Llama family models, we note the following: (i) About 45% of cases show weak or no correlation. (ii) Positive and negative correlations are nearly equally distributed. (iii) The distribution of cases across different correlation degrees is fairly balanced. These results are consistent with observations from other models analyzed (detailed in Appendix B), indicating no definitive correlation between prompt perplexity and model performance.
  • Min-KProb. We also explore the potential relationship between model performance and the prompt’s Min-KProb 31, a method related to pre-training data detection, which judges the model’s familiarity with the text based on the average log-likelihood of the lowest ktokens in the sentence. Intrigued by this, we carry out an analysis akin to that for perplexity, only to find weak positive correlations that lack statistical significance, as shown in the right half of Figure 4.
  • Hidden States. We shift our focus to investigate model perceptions of prompt quality through the lens of model representation, using: (i) Principal Component Analysis (PCA) for direct visualization of hidden states, as shown in Figure 5 (a). The color gradient reflects the model’s performance ranking across 11 prompts per case, revealing no distinct separation between different performance levels. (ii) Probing technique to quantify the extent to which the model’s internal representations are predictive of its performance. We train a reward model (a 3-layer MLP in practice) that scores prompts based on hidden states and apply the standard training loss function 32, which is defined as:
@@ 5 Improving Worst Prompt Performance
Drawing inspiration from the swarm distillation method proposed by Zhou et al. 33, we implement an unsupervised consistency regularization strategy that encourages the model’s predictions for various paraphrases to converge.
@@
    where $\bar{S_{ij}}$ is the mean of $S_{ij}$ over j. The value of $C(X,Y)$ ranges from 0 to $\infty$, reflecting the degree of consistency in the model’s log confidence of generating outputs ($Y$) based on given instructions ($X$), with 0 indicating totally consistent. We obverse that the consistency always decreases at first, followed by increasing, which aligns with the findings in Zhou et al. 33. We train models based on LoRA and stop training at the checkpoint from which $C(X,Y)$ starts to increase.
@@ 6 Conclusion
Appendix A Benchmark Construction
Metric Number of Paraphrases
1 2 3 4 5 6 7 8 9 10 11
Worst 29.18 21.01 17.48 15.40 13.95 12.85 11.95 11.19 10.53 9.93 9.38
Best 29.18 37.34 41.97 45.17 47.55 49.42 50.94 52.18 53.23 54.11 54.86
Best - Worst 0.00 16.33 24.49 29.77 33.60 36.57 38.98 40.99 42.70 44.18 45.48
Table 5: The average score of the Llama-2-70B-chat model across all possible combinations of different numbers of paraphrases.
A.1 Determining the Number of Paraphrases
Basically, the worst performance decreases and the best performance increases monotonically with the number of paraphrases. While it is impractical to list all possible paraphrases and calculate the exact values, we find that these scores converge quickly with a sufficiently large number of paraphrases. Taking Llama-2-70B as an example, we calculate the worst and best performances, along with their difference, for each combination of $n$ (ranging from 1 to 11) paraphrases of a case. Then, we report the average of these scores across all combinations. The results are presented in Table 5. As the results show, with the increase in the number of paraphrases, the average of the model’s best/worst performance initially changes significantly and then levels off. For other models, we observe the same trend of changes. To balance the evaluation efficiency and robustness, we decided to construct 10 paraphrases for each query.
A.2 Instruction for Paraphrasing